# Real enumerative invariants relative to the toric boundary and their refinement

### v1

@@ 6.2 Walls in the space of constraints
Let $[{\Phi}:(\widehat{C},{\boldsymbol{p}})\to{\mathbb{P}}^{2}]\in\overline{\mathcal{M}}_{2,n}^{\mathbb{R}}(\Delta,\widehat{\boldsymbol{w}}(t))_{0\leq t\leq 1}$. In this section, we assume that $\widehat{C}$ does not contain contractible components, i.e., we just contract them (if any) and we call the singularities of $\widehat{C}$ “nodes”. Accordingly, we allow the points of $\widehat{\boldsymbol{p}}$ to collate. Note that such a contraction does not affect the arithmetic genus, since a generic curve $\widehat{C}=S_{2}$ has three real ovals.
@@
    $\widehat{C}$ is a smooth curve of genus $2$, $\widehat{\boldsymbol{w}}_{\partial}(t^{*})$ consists of $n-1$ distinct points, ${\Phi}$ is birational onto its image, but not an immersion; furthermore, the germs ${\Phi}(\widehat{C},p_{i})$, $i=1,2$, and the germs ${\Phi}(\widehat{C},p_{i}^{(s)})$ for all $s=0,1,2$ and $k_{i}^{(s)}\geq 2$ are smooth;
@@
    $\widehat{C}$ is a smooth curve of genus $2$, $\widehat{\boldsymbol{w}}_{\partial}$ contains $n-2$ distinct points so that $w_{i}^{(s)}(t^{*})=w_{j}^{(s)}(t^{*})$ for some $s=0,1,2$ and $i\neq j$, and $p_{i}^{(s)}=p_{j}^{(s)}=p\in{\boldsymbol{p}}$; in addition, the map ${\Phi}$ is an immersion;
@@
    $\widehat{C}=\widehat{C}_{1}\cup\widehat{C}_{2}$, where $g(\widehat{C}_{1})+g(\widehat{C}_{2})=2$, and the only node joining $\widehat{C}_{1}$ and $\widehat{C}_{2}$ is mapped to (×$({\mathbb{R}}^{\times})^{2}\subset{\mathbb{P}}_{\mathbb{R}}^{2}$; $\widehat{\boldsymbol{w}}_{\partial}(t^{*})$ consists of $n-1$ distinct points; the maps ${\Phi}:\widehat{C}_{i}\to{\mathbb{P}}^{2}$, $i=1,2$, are birational immersions;
@@
    $\widehat{C}=\widehat{C}_{1}\cup\widehat{C}_{2}$, where $g(\widehat{C}_{1})+g(\widehat{C}_{2})=2$; $\widehat{\boldsymbol{w}}_{\partial}(t^{*})$ consists of $n-1$ distinct points; the only node $p$ joining $\widehat{C}_{1}$ and $\widehat{C}_{2}$ is mapped to some point $w_{i}^{(s)}(t^{*})\in\widehat{\boldsymbol{w}}(t^{*})$; the maps ${\Phi}:\widehat{C}_{i}\to{\mathbb{P}}^{2}$, $i=1,2$, are birational immersions; the germ ${\Phi}(\widehat{C}_{j},p)$ intersects $D_{s}$ with multiplicity $2l_{j}$, $j=1,2$, where $l_{1}+l_{2}=k_{i}^{(s)}$ and these two germs intersect each other with multiplicity $\min\{2l_{1},2l_{2}\}$;
@@
Proof. (1i) We can assume that the sequences $\{{\boldsymbol{w}}(t)\}_{0\leq t\leq 1}$ avoid $\tau_{0}$-neighborhoods of the intersection points of the coordinate axes, for some $\tau_{0}>0$. Assume that ${\Phi}(\widehat{C})$ contains some coordinate axis.
Observe that if ${\Phi}(\widehat{C})$ contains at least one coordinate axis, then it contain all of them. Indeed, otherwise we would have ${\boldsymbol{w}}(t)$ arbitrarily close to the intersection points of the coordinate axes, contrary to the above assumption. Furthermore, the intersection points of the coordinate axes must lift to nodes of $\widehat{C}$, and they smooth out along the path $\{{\boldsymbol{w}}(t)\}_{0\leq t\leq 1}$. In particular, the union $\widehat{C}_{1}$ of the components of $\widehat{C}$ mapped onto the coordinate axes is a curve of arithmetic genus $\geq 1$.
The union $\widehat{C}_{2}$ of the other components of $\widehat{C}$ cannot be mapped onto a plane curve of odd degree. Indeed, otherwise, one would get an odd number of real branches of ${\Phi}(\widehat{C}_{2})$ intersecting $D_{0}$ with odd multiplicity yielding a one-dimensional real arc in the quadrant $\{x_{1}/x_{0}<0,\ x_{2}/x_{0}<0\}$ which is not possible. Thus, ${\Phi}_{*}(\widehat{C}_{2})$ has an even degree. If $\widehat{C}_{2}$ has an arithmetic genus 00, then the real part of ${\Phi}(\widehat{C}_{2})$ contains a connected graph passing through $w_{1}$ and $w_{2}$, and then there should be four real local branches of ${\Phi}(\widehat{C}_{2})$ centered on $D_{1,+}\cup D_{2,+}$ that intersect $D_{1}\cup D_{2}$ with odd multiplicity. Hence, the centers of these branches must lift to the nodes of $\widehat{C}$, which would yield that the arithmetic genus of $\widehat{C}$ is at least $4$ a contradiction. Thus, $\widehat{C}_{1}$ and $\widehat{C}_{2}$ both have arithmetic genus $1$, and are joined by one node. It is easy to see, that this node must be mapped by ${\Phi}$ to the center of the unique local branch of ${\Phi}(\widehat{C}_{2})$ on $D_{0,+}$. This local branch must lie in the closure of $({\mathbb{P}}_{\mathbb{R}}^{2})^{+}$. Hence the image of one connected component of the real part of $\widehat{C}_{2}$ must pass through, say, $w_{1}$ and contain the real local branch centered on $D_{0,+}$. It follows that ${\Phi}(\widehat{C}_{2})$ has two real local branches centered on $D_{1,=}$ and intersecting $D_{1}$ with odd multiplicity. Thus, their centers must lift to two nodes of $\widehat{C}$ joining $\widehat{C}_{1}$ and $\widehat{C}_{2}$ making the arithmetic genus of $\widehat{C}$ at least $3$ which is a contradiction.
@@
Third, $\widehat{C}$ cannot be the union of a rational and an elliptic component. Then we obtain an extra Menelaus condition on ${\boldsymbol{w}}_{\partial}$, while the parts of ${\boldsymbol{w}}_{\partial}$ lying on the image of the rational and of the elliptic components are in general position. Thus, ${\boldsymbol{w}}_{\partial}$ and one of the points $w_{1},w_{2}$ determine a finite set of pairs of a rational and an elliptic components of $\widehat{C}$ (Lemma 2.4), which imposes a condition on the location of the second point in $\{w_{1},w_{2}\}$, hence, defining a stratum of codimension $\geq 2$ in ${\mathcal{W}}$.
(1iii) Suppose that $\widehat{C}$ is irreducible, ${\Phi}:\widehat{C}\twoheadrightarrow C\hookrightarrow{\mathbb{P}}^{2}$ a multiple covering. By the Riemann-Hurwitz formula, $C$ cannot be of geometric genus $2$. Also, $C$ cannot be rational since otherwise, the fixed position of ${\boldsymbol{w}}_{\partial}$ determines a finite set of rational curves, and hence two conditions are imposed on the position of $w_{1},w_{2}$. This defines a stratum of codimension $\geq 2$ in ${\mathcal{W}}$ avoided by the path $\{{\boldsymbol{w}}(t)\}_{0\leq t\leq 1}$. The only remaining option is an elliptic curve $C$. As in the case of the rational $C$, we derive that ${\boldsymbol{w}}_{\partial}$ varies in an open dense subset of ${\mathcal{W}}_{\partial}$, while the position of $w_{1},w_{2}$ is constrained by one condition. It follows that all the points of the sequence $\widehat{\boldsymbol{w}}_{\partial}$ are distinct, and hence they all must be ramification points of the covering. However, #$\#\widehat{\boldsymbol{w}}_{\partial}\geq 3$ which yields a contradiction in the Riemann-Hurwitz formula.
In view of (1ii), it remains to study the case of two-component $\widehat{C}=\widehat{C}_{1}\cup\widehat{C}_{2}$ such that either $g(\widehat{C}_{1})=2$, $g(\widehat{C}_{2}=0$, or $g(\widehat{C}_{1})=\widehat{C}_{2})=1$. It follows that ${\boldsymbol{w}}_{\partial}$ is constrained by an extra Menelaus condition, and this is the only condition imposed. Hence, all the points of $\widehat{\boldsymbol{w}}_{\partial}$ differ from each other, and the genus of the image of any component must be equal to the genus of that component. The argument of the preceding paragraph leaves the only option of a multiple covering ${\Phi}:\widehat{C}_{2}\twoheadrightarrow C_{2}\hookrightarrow{\mathbb{P}}^{2}$ with rational $\widehat{C}_{2}$ and $C_{2}$. If $m$ is the degree of the covering, then each point $w_{i}^{(s)}\in C_{2}$ is a ramification point of the covering of index $m-1$, and hence we get a contradiction in the Riemann-Hurwitz formula:
@@
(2) The classification of elements of
¯$\overline{\mathcal{M}}_{2,n}(\Delta,\widehat{\boldsymbol{w}}(t^{*}))\setminus{\mathcal{M}}_{2,n}(\Delta,\widehat{\boldsymbol{w}}(t^{*}))$
is derived literally in the same manner as Lemma 5.2(2ii,2iii,2iv,2v0). We only notice in addition that the cases including multiple coverings do not occur as we showed in part (1iii). $\Box$

### v2

@@ 6.2 Walls in the space of constraints
Let $[{\Phi}:(\widehat{C},{\boldsymbol{p}})\to{\mathbb{P}}^{2}]\in\overline{\mathcal{M}}_{2,n}^{\mathbb{R}}(\Delta,\widehat{\boldsymbol{w}}(t))_{0\leq t\leq 1}$. In this section, we assume that $\widehat{C}$ does not contain contractible components, i.e., we just contract them (if any) and we call the singularities of $\widehat{C}$ “nodes”. Note that such a contraction does not affect the arithmetic genus, since a generic curve $\widehat{C}$ has three real ovals. We also allow the points of $\widehat{\boldsymbol{p}}$ to collate. We use the notations of Section 5.2: namely, $S$ denotes a locus in the space of constraints $\operatorname{Men}^{\rho}_{\mathbb{R}}(\Delta)\times({\mathbb{P}}^{2}_{\mathbb{R}})^{(+,-)}\times({\mathbb{P}}^{2}_{\mathbb{R}})^{(-,+)}$ and ${\operatorname{pr}}(S)$ denotes the image of ¯$\overline{\mathcal{M}}_{2,n}^{\mathbb{R}}(\Delta,\widehat{\boldsymbol{w}})$ in the linear system $|{\mathcal{L}}_{P_{\Delta}}|$ when $\widehat{\boldsymbol{w}}$ ranges over $S$.
In view of the results of Section 6.1, we can ignore the walls along which ¯$\overline{\mathcal{M}}^{\mathbb{R}}_{2,n}(\Delta,\widehat{\boldsymbol{w}})={\mathcal{M}}^{\mathbb{R}}_{2,n}(\Delta,\widehat{\boldsymbol{w}})$. The following lemma classifies other walls.
@@
(2a) either exactly one of the points $w_{1}(t^{*})$, $w_{2}(t^{*})$ is a singular point of the curve $C={\Phi}(\widehat{C})$, and the following holds:
@@
    $\widehat{\boldsymbol{w}}_{\partial}(t^{*})$ consists of $n$ distinct points, the curve $\widehat{C}$ is smooth of genus $2$, ${\Phi}$ is an immersion onto a curve $C$ which is smooth along the toric divisors, while $w_{i}(t^{*})$ is a center of at least one real local branch, $i=1,2$;
(2b) or $w_{1}(t^{*})$, $w_{2}(t^{*})$ are smooth points of the curve $C={\Phi}(\widehat{C})$, and one of the following holds:
@@
    $\widehat{C}$ is a smooth curve of genus $2$, $\widehat{\boldsymbol{w}}_{\partial}(t^{*})$ consists of $n-1$ distinct points, ${\Phi}$ is birational onto its image, but not an immersion: it may have singular local branches centered in ${\operatorname{Tor}}(P_{\Delta})^{\times}$ as well as singularities of type $A_{2m}$ at some points $w_{i}^{\sigma}\in\widehat{\boldsymbol{w}}_{\partial}(t^{*})$, where $m\geq k_{i}^{\sigma}$;
@@
    $\widehat{C}$ is a smooth curve of genus $2$, $\widehat{\boldsymbol{w}}_{\partial}$ contains $n-2$ distinct points so that $w_{i}^{(s)}(t^{*})=w_{j}^{(s)}(t^{*})=:w$ for some $s\in\{0,1,2\}$ and $i\neq j$; the map ${\Phi}$ is birational onto its image $C$, where the curve $C$ is immersed outside $\widehat{\boldsymbol{w}}_{\partial}(t^{*})\setminus\{w\}$ and may have singularities of type $A_{2m}$ at some points of $\widehat{\boldsymbol{w}}_{\partial}(t^{*})\setminus\{w\}$, while at $w$, $C$ either is smooth or has two smooth branches;
@@
    $\widehat{C}=\widehat{C}_{1}\cup\widehat{C}_{2}$, where $g(\widehat{C}_{1})+g(\widehat{C}_{2})=2$, $\widehat{\boldsymbol{w}}_{\partial}(t^{*})$ consists of $n-1$ distinct points, and the maps ${\Phi}:\widehat{C}_{i}\to{\mathbb{P}}^{2}$, $i=1,2$, are immersions; furthermore, the only node $p$ joining $\widehat{C}_{1}$ and $\widehat{C}_{2}$ either is mapped to (×$({\mathbb{R}}^{\times})^{2}\subset{\mathbb{P}}_{\mathbb{R}}^{2}$, or is mapped to some point $w_{i}^{(s)}(t^{*})\in\widehat{\boldsymbol{w}}(t^{*})$; the germ ${\Phi}(\widehat{C}_{j},p)$ intersects $D_{s}$ with multiplicity $2l_{j}$, $j=1,2$, where $l_{1}+l_{2}=k_{i}^{(s)}$ and these two germs intersect each other with multiplicity $\min\{2l_{1},2l_{2}\}$;
@@
Proof. (1i) We can assume that the sequences $\{{\boldsymbol{w}}(t)\}_{0\leq t\leq 1}$ avoid $\rho_{0}$-neighborhoods of the intersection points of the coordinate axes, for some $\rho_{0}>0$.
Suppose that ${\Phi}(\widehat{C})$ contains some coordinate axis. In this case, ${\Phi}(\widehat{C})$ contains all of them. Indeed, otherwise, we would have ${\boldsymbol{w}}(t)$ arbitrarily close to the intersection points of the coordinate axes, contrary to the above assumption. Furthermore, the intersection points of the coordinate axes must lift to nodes of $\widehat{C}$, and they smooth out along the path $\{{\boldsymbol{w}}(t)\}_{0\leq t\leq 1}$. In particular, the union $\widehat{C}_{1}$ of the components of $\widehat{C}$ mapped onto the coordinate axes is a curve of arithmetic genus $\geq 1$.
The union $\widehat{C}_{2}$ of the other components of $\widehat{C}$ cannot be mapped onto a plane curve of odd degree. Indeed, otherwise, one would get an odd number of real branches of ${\Phi}(\widehat{C}_{2})$ intersecting $D_{0}$ with odd multiplicity yielding a one-dimensional real arc in the quadrant $\{x_{1}/x_{0}<0,\ x_{2}/x_{0}<0\}$ which is not possible. Thus, ${\Phi}_{*}(\widehat{C}_{2})$ has an even degree. If $\widehat{C}_{2}$ has an arithmetic genus 00, then the real part of ${\Phi}(\widehat{C}_{2})$ contains a connected graph passing through $w_{1}$ and $w_{2}$, and then there should be four real local branches of ${\Phi}(\widehat{C}_{2})$ centered on $D_{1}^{+}\cup D_{2}^{+}$ (see Section 3.4 for notations) that intersect $D_{1}\cup D_{2}$ with odd multiplicity. Hence, the centers of these branches must lift to the nodes of $\widehat{C}$, which would yield that the arithmetic genus of $\widehat{C}$ is at least $4$; a contradiction. Thus, $\widehat{C}_{1}$ and $\widehat{C}_{2}$ both have arithmetic genus $1$, and are joined by one node. It is easy to see, that this node must be mapped by ${\Phi}$ to the center of the unique local branch of ${\Phi}(\widehat{C}_{2})$ on $D_{0}^{+}$. This local branch must lie in the closure of $({\mathbb{P}}_{\mathbb{R}}^{2})^{+}$. Hence the image of one connected component of the real part of $\widehat{C}_{2}$ must pass through, say, $w_{1}$ and contain the real local branch centered on $D_{0}^{+}$. It follows that ${\Phi}(\widehat{C}_{2})$ has two real local branches centered on $D_{1}^{+}$ and intersecting $D_{1}$ with odd multiplicity. Thus, their centers must lift to two nodes of $\widehat{C}$ joining $\widehat{C}_{1}$ and $\widehat{C}_{2}$ making the arithmetic genus of $\widehat{C}$ at least $3$ which is a contradiction.
@@
Third, $\widehat{C}$ cannot be the union of a rational and an elliptic component. Then, we obtain an extra Menelaus condition on ${\boldsymbol{w}}_{\partial}$, while the parts of ${\boldsymbol{w}}_{\partial}$ lying on the image of the rational and the elliptic components are in general position. Thus, ${\boldsymbol{w}}_{\partial}$ and one of the points $w_{1},w_{2}$ determine a finite set of pairs of a rational and an elliptic components of $\widehat{C}$ (Lemma 2.4), which imposes a condition on the location of the second point in $\{w_{1},w_{2}\}$, hence, defining a stratum of codimension $\geq 2$ in ${\mathcal{W}}$.
(1iii) Suppose that $\widehat{C}$ is irreducible and ${\Phi}:\widehat{C}\twoheadrightarrow C\hookrightarrow{\mathbb{P}}^{2}$ is a multiple covering. By the Riemann-Hurwitz formula, the curve $C$ cannot be of geometric genus $2$. Also, $C$ cannot be rational since otherwise, the fixed position of ${\boldsymbol{w}}_{\partial}$ determines a finite set of rational curves, and hence two conditions are imposed on the position of $w_{1},w_{2}$. This defines a stratum of codimension $\geq 2$ in ${\mathcal{W}}$ avoided by the path $\{{\boldsymbol{w}}(t)\}_{0\leq t\leq 1}$. The only remaining option is an elliptic curve $C$. As in the case of the rational curve $C$, we derive that ${\boldsymbol{w}}_{\partial}$ varies in an open dense subset of ${\mathcal{W}}_{\partial}$, while the position of $w_{1},w_{2}$ is constrained by one condition. It follows that all the points of the sequence $\widehat{\boldsymbol{w}}_{\partial}$ are distinct, and hence they all must be ramification points of the covering. However, #$\#\widehat{\boldsymbol{w}}_{\partial}\geq 3$ which yields a contradiction in the Riemann-Hurwitz formula.
In view of (1ii), it remains to study the case of two-component $\widehat{C}=\widehat{C}_{1}\cup\widehat{C}_{2}$ such that either $g(\widehat{C}_{1})=2$, $g(\widehat{C}_{2})=0$, or $g(\widehat{C}_{1})=g(\widehat{C}_{2})=1$. It follows that ${\boldsymbol{w}}_{\partial}$ is constrained by an extra Menelaus condition, and this is the only condition imposed. Hence, all the points of $\widehat{\boldsymbol{w}}_{\partial}$ differ from each other, and the genus of the image of any component must be equal to the genus of that component. The argument of the preceding paragraph leaves the only option of a multiple covering ${\Phi}:\widehat{C}_{2}\twoheadrightarrow C_{2}\hookrightarrow{\mathbb{P}}^{2}$ with rational $\widehat{C}_{2}$ and $C_{2}$. If $m$ is the degree of the covering, then each point $w_{i}^{(s)}\in C_{2}$ is a ramification point of the covering of index $m-1$, and hence we get a contradiction in the Riemann-Hurwitz formula:
@@
(2) We mainly follow the argument in parts (2ii)-(2v) of the proof of Lemma 5.2. The items (2i), (2iv), (2v) of the present lemma can be treated in the same manner as the items (2i), (2iv), (2v) in Lemma 5.2. The case of (2v) appears to be even much simpler, since multiple coverings are not possible as we established in the preceding part of the proof. The cases (2ii) and (2iii) require additional arguments.
(2ii) Suppose that $\widehat{\boldsymbol{w}}_{\partial}(t^{*})$ consists of $n$ distinct points, $\widehat{C}$ is a smooth curve of genus $2$, ${\Phi}:\widehat{C}\to C={\Phi}(\widehat{C})$ is birational, but not an immersion. To verify the conditions of item (2ii) of the lemma, it is enough to consider the case of $C$ having a singular branch at some point $w_{i}^{\sigma}\in\widehat{\boldsymbol{w}}_{\partial}(t^{*})$. Fixing the position of the point $w_{i}^{\sigma}$, we obtain a subfamily ${\operatorname{pr}}(S^{fix}(w_{i}^{\sigma}))\subset{\operatorname{pr}}(S)$ of dimension at least $n-2\geq 3$ and then apply [16, Inequality (5) in Lemma 2.1] (cf. (31)):
$-CK_{{\operatorname{Tor}}(P_{\Delta})}\geq-2+\left(-CK_{{\operatorname{Tor}}(P_{\Delta})}-n+2\right)+(n-3)+({\operatorname{mult}}Q-1)$
$=-CK_{{\operatorname{Tor}}(P_{\Delta})}-3+({\operatorname{mult}}Q-1)\quad\Longrightarrow\quad{\operatorname{mult}}Q\leq 4\ ,$
where $Q$ is the local branch of $C$ at $w_{i}^{\sigma}$. It remains to show that ${\operatorname{mult}}Q=2$. Arguing by contradiction, assume that ${\operatorname{mult}}Q=3$ or $4$. Without loss of generality, we can assume that the germ of ${\operatorname{pr}}(S)\subset|{\mathcal{L}}_{P_{\Delta}}|$ at $C$ is an equisingular family. According to items (a) and (b) in Section 2.3, we obtain the following upper bound completely analogous to (5):
$\dim{\operatorname{pr}}(S^{fix}(w_{i}^{\sigma}))\leq h^{0}(C,{\mathcal{J}}_{Z/C}\otimes_{{\mathcal{O}}_{C}}{\operatorname{in}}^{*}{\mathcal{L}}_{P_{\Delta}})=h^{0}\left(\widehat{C},{\mathcal{O}}_{\widehat{C}}(-{\boldsymbol{d}})\otimes_{{\mathcal{O}}_{\widehat{C}}}{\Phi}^{*}{\mathcal{L}}_{P_{\Delta}}\right),$ (44)
where
$\deg{\boldsymbol{d}}=2\sum_{z\in{\operatorname{Sing}}\hskip 1.70709pt(C)}\delta(C,z)+(-CK_{{\operatorname{Tor}}(P_{\Delta})}-n+2)+r,\quad 2\leq r\leq 3,$
$\Longrightarrow\quad\deg({\mathcal{O}}_{\widehat{C}}(-{\boldsymbol{d}})\otimes_{{\mathcal{O}}_{\widehat{C}}}{\Phi}^{*}{\mathcal{L}}_{P_{\Delta}})=C^{2}-\deg{\boldsymbol{d}}=n-r\geq\begin{cases}3,\ &r=2,\\ 2,\ &r=3\end{cases}.$
In case of $r=2$, we have $h^{1}({\mathcal{O}}_{\widehat{C}}(-{\boldsymbol{d}})\otimes_{{\mathcal{O}}_{\widehat{C}}}{\Phi}^{*}{\mathcal{L}}_{P_{\Delta}})=0$ since $22g(\widehat{C})-2=$ 2. Thus, by the Riemann-Roch formula, we have
$\dim{\operatorname{pr}}(S^{fix}(w_{i}^{\sigma}))\leq h^{0}({\mathcal{O}}_{\widehat{C}}(-{\boldsymbol{d}})\otimes_{{\mathcal{O}}_{\widehat{C}}}{\Phi}^{*}{\mathcal{L}}_{P_{\Delta}})=n-3$
contradicting the lower bound $\dim{\operatorname{pr}}(S^{fix}(w_{i}^{\sigma}))\geq n-2$ pointed above. In case of $r=3$, we have
$h^{0}({\mathcal{O}}_{\widehat{C}}(-{\boldsymbol{d}})\otimes_{{\mathcal{O}}_{\widehat{C}}}{\Phi}^{*}{\mathcal{L}}_{P_{\Delta}})\leq h^{8}({\mathcal{O}}_{\widehat{C}}(-{\boldsymbol{d}}^{\prime})\otimes_{{\mathcal{O}}_{\widehat{C}}}{\Phi}^{*}{\mathcal{L}}_{P_{\Delta}}),$
where ${\boldsymbol{d}}^{\prime}={\boldsymbol{d}}-w$ with $w$ being one of the points of ${\boldsymbol{d}}$. Thus,
$\deg{\boldsymbol{d}}^{\prime}=\deg{\boldsymbol{d}}-1\quad\Longrightarrow\quad\deg({\mathcal{O}}_{\widehat{C}}(-{\boldsymbol{d}}^{\prime})\otimes_{{\mathcal{O}}_{\widehat{C}}}{\Phi}^{*}{\mathcal{L}}_{P_{\Delta}})=n-r+1\geq 3,$
and we derive the same contradiction.
(2iii) Proceeding along the lines of part (2iii) of the proof of Lemma 5.2, we need to verify suitable analogs of the estimates (33), (34), and (35) in the genus $2$ case.
An analog of (33) reads
$-CK_{{\operatorname{Tor}}(P_{\Delta})}\geq-2+\left(-CK_{{\operatorname{Tor}}(P_{\Delta})}-n+4\right)+(n-4)+\sum_{B}({\operatorname{mult}}Q-1)$
$=-CK_{{\operatorname{Tor}}(P_{\Delta})}-2+\sum_{B}({\operatorname{mult}}Q-1)\quad\Longrightarrow\quad\sum_{B}({\operatorname{mult}}Q-1)\leq 2\ ,$
where $B$ is the set of all singular local branches of $C$ in ${\operatorname{Tor}}(P_{\Delta})^{\times}\cup\{w,w^{\prime}\}$ (recall that $w^{\prime}$ is any point in $\widehat{\boldsymbol{w}}_{\partial}(t^{*})\setminus\{w\}$). We have analogs of (34) and (35):
$\dim{\operatorname{pr}}(S^{fix}(w))\leq h^{0}(\widehat{C},{\mathcal{O}}_{\widehat{C}}(-{\boldsymbol{d}}_{1})\otimes_{{\mathcal{O}}_{\widehat{C}}}{\Phi}^{*}{\mathcal{L}}_{P_{\Delta}}),$
where
$\deg{\boldsymbol{d}}_{1}=2\sum_{z\in{\operatorname{Sing}}\hskip 1.70709pt(C)}\delta(C,z)+(-CK_{{\operatorname{Tor}}(P_{\Delta})}-n+3)+r,\quad r=1\ \text{or}\ 2,$
and hence
$\deg(\widehat{C},{\mathcal{O}}_{\widehat{C}}(-{\boldsymbol{d}}_{1})\otimes_{{\mathcal{O}}_{\widehat{C}}}{\Phi}^{*}{\mathcal{L}}_{P_{\Delta}})=C^{2}-\deg{\boldsymbol{d}}_{1}=C^{2}-(C^{2}+CK_{{\operatorname{Tor}}(P_{\Delta})}-2)$
$-(-CK_{{\operatorname{Tor}}(P_{\Delta})}-n+3)-r=n-1-r\geq\begin{cases}3,\ &r=1,\\ 2,\ &r=2.\end{cases}$
Then, we literally follow the last paragraph of part (2ii) of the proof of the lemma and derive that $\dim{\operatorname{pr}}(S^{fix}(w))\leq n-3$ contrary to the initial lower bound $\dim{\operatorname{pr}}(S^{fix}(w))\geq n-2$. This contradiction completes the study of the case (2iii) of the lemma. $\Box$